# Evaluating High-Resolution Piano Sustain Pedal Depth Estimation with Musically Informed Metrics

### v1

@@ Abstract
Evaluation for continuous piano pedal depth estimation tasks remains incomplete when relying only on conventional frame-level metrics, which overlook musically important features such as direction-change boundaries and pedal curve contours. To provide more interpretable and musically meaningful insights, we propose an evaluation framework that augments standard frame-level metrics with an action-level assessment measuring direction and timing using segments of press/hold/release states and a gesture-level analysis that evaluates contour similarity of each press-release cycle. We apply this framework to compare an audio-only baseline with two variants: one incorporating symbolic information from MIDI, and another trained in a binary-valued setting, all within a unified architecture. Results show that the MIDI-informed model significantly outperforms the others at action and gesture levels, despite modest frame-level gains. These findings demonstrate that our framework captures musically relevant improvements indiscernible by traditional metrics, offering a more practical and effective approach to evaluating pedal depth estimation models.
@@ 1 Introduction
Although the sustain pedal is a continuous control crucial for producing distinct sound effects essential to classical piano playing, previous studies treated it as an auxiliary signal within piano transcription systems [1, 2, 3, 4] and mostly focused on binary on/off detection [5, 6, 7, 8, 9, 10]. Previous work [11] demonstrates the importance and benefits of continuous pedal depth estimation. However, compared with mature MIR tasks such as melody extraction, beat tracking, and chord recognition, pedal depth estimation remains underexplored, and its evaluation still largely relies on standard generic frame-level metrics (MSE, MAE, F1, etc.), which are not specifically designed for musically meaningful information, such as whether action boundaries are correct, and whether phrase-level contours are reasonable. Binary F1-score can ignore important cases such as half-pedal and pedal fluttering, while multi-class F1-score overlooks expressive musical aspects in the curve. Frame-level regression metrics (MSE/MAE) better reflect continuous values, but still over-penalize small timing shifts or minor jitters, even when two pedal curves are nearly equivalent in perceived effect.
@@
In this paper, we propose evaluating pedal depth estimation models not only with existing frame-level metrics but also using two additional levels of metrics that are musically more interpretable and practically more informative: (1) an action-level evaluation, which assesses predictions based on segments of press/hold/release; and (2) a gesture-level evaluation, which treats each press–release cycle as a unit and compares contour similarity within that unit. Evaluation at action and gesture levels extends the generic frame-based analysis by providing richer insights into model performance that align with actual practice and pedagogical concepts in piano playing.
For our experiments, we design three models within a unified Transformer-based architecture to demonstrate how the three levels yield complementary, interpretable judgments. The goal is to explore whether our new framework can reveal differences in model performance missed in frame-level metrics, and whether evaluation at the action or gesture level will help us better understand model behavior in musical contexts. Results show that continuous-valued models outperform binary-valued at predicting all press/release actions and plotting gesture shapes; adding MIDI further strengthens action boundary clarity and gesture contour matching, though fine micro-dynamics in long mountain-like gestures remain challenging for our models. Overall, this framework provides more musically meaningful analysis and thus more informative model performance diagnosis for piano sustain pedal depth estimation tasks.
@@ 3.1 Pedal Prediction Model Variants
    Binary: binarized labels trained with BCE; output in $[0,1]$.
@@
    Audio: continuous regression on mel bins + MFCCs.
@@ 3.2 Frame-Level Evaluation
Standardized evaluation protocols for continuous pedal estimation are still under discussion; most evaluations still borrow criteria from binary prediction [11]. Conventional frame-wise metrics measures: (i) classification scores after discretization (binary and 4-class) using Precision/Recall/F1, and (ii) regression errors on the continuous depth $x_{1:T}$ using MAE and MSE. These summarize per-frame signal fidelity and enable comparison to prior work.
@@ 4 Evaluation Framework
Beyond frame-level metrics, we propose two novel perspectives for evaluation inspired by piano pedagogy: (i) an action-based evaluation (Section 4.1), in which we segment the pedal curve into press/hold/release and score temporal coverage and alignment; and (ii) a gesture-based evaluation (Section 4.2), where we consider complete pedal gestures from pedal on to off and compare characteristic contour shapes that implies pedaling in expressive piano performance.
@@ 4.1 Action-Based Evaluation
The action-based evaluation can now be framed as a three-class classification task for press, hold, release classes using action states extracted from by the steps mentioned above, and the evaluation follows the standard precision, recall, and F1 framework. We report scores for each action class as well as class-balanced macro and weighted averages, all shown in Table 2.
[…]
[11] 0.8975 0.8971 0.8973 0.6849 0.6971 0.6863 0.0425 0.1339
Binary 0.8929 0.8909 0.8914 0.5622 0.6615 0.5655 0.0880 0.1714
Audio 0.9043 0.9037 0.9039 0.7013 0.7153 0.7045 0.0416 0.1237
Audio+MIDI 0.9379 0.9370 0.9372 0.7529 0.7661 0.7546 0.0280 0.0986
Table 1: Frame-level results.
@@ 4.2 Gesture-Based Evaluation
For all non-plain gestures in the corpus: duration (frames) has mean $147.8$, median $70.0$, std $402.9$; max-depth ratio has mean $0.645$, median $0.688$, std $0.220$. We set our max-depth ratio threshold to $\theta=0.65$ and duration threshold to $100$ frames to define our quadrants.
We already see that frame-wise scores can miss contour similarity in Section 3.3. To compare gesture shapes while tolerating small local deviations, we adopt (i) Fourier-descriptor analysis: compute discrete Fourier coefficients and reconstruct with the first K=11K{=}11 coefficients to suppress high-frequency noise/model artifacts, and compare the reconstructed low-pass signals via MSE; (ii) 5-point analysis: compute duration-weighted MSE using only five landmarks per gesture (start, end, median, mean, max). According to these new metrics, the top example in Figure 2 yields an MSE of 0.0586 using the Fourier method, whereas the bottom example obtains an MSE of 0.0274 in the 5-point analysis, now both being less harshly graded than by raw MSE scores.
@@ 5.2 Evaluation Across Models
We use conventional frame-based metrics (Table 1) as well as our action- (Table 2) and gesture-based metrics (Table 3). Overall, the Audio+MIDI model performs the best, followed by Audio, with Binary the last. This pattern highlights that estimating continuous pedal values is beneficial, and adding MIDI input further enhances performance by providing helpful structural priors.
Continuous estimation matters (vs. binary classification). Although Binary achieves a comparable frame-wise F1 (0.8914), it exhibits systematic bias at both action and gesture levels (Figure 3, Table 2). At action level, recall scores for press and release are very low, indicating that many actions are missed. This likely comes from data imbalance in binary setting where most frames are “off,” encouraging under-prediction of pedal activation. At gesture level, Binary largely collapses to predicting highland and pinnacle, and rarely recognizes more complex patterns such as mountain (Fig. 3). Even for highland and pinnacle, which Binary predicts frequently, gesture-level MSE remains unsatisfactory (Table 3).
Models trained for continuous estimation have remarkably improved action-level recalls and modestly but consistently higher precisions than Binary (Table 2). In addition, at the gesture level, Audio and Audio+MIDI can identify gesture categories other than highland and pinnacle, yielding distributions closer to ground truth (Figure 3), and they also result in substantially lower per-gesture MSE, a positive sign for more accurate shape modeling. These findings agree with [11] and reinforce the importance of continuous-valued prediction for piano pedal estimation.
@@ 6 Conclusion
We studied sustain pedal depth estimation with controlled experiments and evaluated using our proposed new framework with three levels at frame, action, and gesture. Across three model variants, we demonstrated that continuous-valued estimation is crucial for accurately capturing actions and preserving gesture characteristics and pre-extracted MIDI pitch velocity information provides useful structural priors for better contour shaping in gestures. While conventional frame-wise metrics can miss musically acceptable behavior, our proposed three-level approach provides a more comprehensive view of strengths and limitations of our models. These results shows the value of our musically informed evaluation and point to future work on boundary-sensitive objectives, contour-aware losses, and perceptual validation for piano sustain pedal depth estimation tasks.

### v2

@@ Abstract
Evaluation of continuous piano pedal depth estimation tasks remains incomplete when relying only on conventional frame-level metrics, which overlook musically important features such as direction-change boundaries and pedal curve contours. To provide more interpretable and musically meaningful insights, we propose an evaluation framework that augments standard frame-level metrics with an action-level assessment measuring direction and timing using segments of press/hold/release states and a gesture-level analysis that evaluates contour similarity of each pedal press-release cycle. We apply this framework to compare an audio-only baseline with two variants: one incorporating symbolic information from MIDI, and another trained on binary-valued targets (pedal on/off only), all within a unified architecture. Results show that the MIDI-informed model significantly outperforms the others at action and gesture levels, despite modest frame-level gains. These findings demonstrate that our framework captures musically relevant improvements indiscernible by traditional metrics, offering a more practical and effective approach to evaluating pedal depth estimation models.
@@ 1 Introduction
Although the sustain pedal is a continuous control crucial for producing distinct sound effects essential to classical piano playing, previous studies treated it as an auxiliary signal within piano transcription systems [7, 9, 25, 26] and mostly focused on binary on/off detection [12, 11, 13, 14, 15, 16]. Previous work [5] demonstrates the importance and benefits of continuous pedal depth estimation. However, compared with mature MIR tasks such as melody extraction, beat tracking, and chord recognition, pedal depth estimation remains underexplored, and its evaluation still largely relies on standard generic frame-level metrics (MSE, MAE, F1, etc.), which are not specifically designed for musically meaningful information, such as whether action boundaries are correct and whether phrase-level contours are reasonable. Binary F1-score can ignore important cases such as half-pedal and pedal fluttering, while multi-class F1-score overlooks expressive musical aspects of the pedal curve. Frame-level regression metrics (MSE/MAE) better reflect continuous values, but still over-penalize small timing shifts or minor jitters, even when two pedal curves are nearly equivalent in perceived musical effect.
@@
In this paper, we propose evaluating pedal depth estimation models not only with existing frame-level metrics but also using two additional levels of metrics that are musically more interpretable and practically more informative: (1) an action-level evaluation, which assesses predictions based on press/hold/release segments; and (2) a gesture-level evaluation, which treats each press–release cycle as a unit and compares contour similarity within that unit. Evaluation at action and gesture levels extends the generic frame-based analysis by providing richer insights into model performance that align with actual practice and pedagogical concepts in piano playing.
To validate our proposed evaluation framework, we compare three model variants with expected capability gaps: audio-only (baseline), MIDI-informed, and audio-only but trained on binarized (on/off) targets, all built on the same Transformer-based architecture. The goal is to test whether our framework can reveal performance differences that frame-level metrics fail to capture, and whether action- and gesture-level evaluation can better characterize model behavior in musical contexts. Results show that the binary-target model underperforms at predicting action states and capturing gesture shapes; adding MIDI to baseline further strengthens action boundary clarity and gesture contour matching, though fine micro-dynamics in long mountain-like gestures remain challenging for our models. Overall, this framework provides more musically meaningful analysis and thus more informative diagnosis of model performance for piano sustain pedal depth estimation tasks.
@@ 3.1 Pedal Prediction Model Variants
    Audio (Binary): binarized labels trained with BCE; raw sigmoid output (before thresholding) used as predicted depth.
@@
    Audio: baseline continuous regression on audio features (mel bins + MFCCs).
@@ 3.2 Frame-Level Evaluation
Standardized evaluation protocols for continuous pedal estimation are still under discussion; most evaluations still borrow criteria from binary prediction [5]. Conventional frame-wise metrics measure: (i) classification scores after discretization (binary and 4-class) using Precision/Recall/F1, and (ii) regression errors on the continuous depth $x_{1:T}$ using MAE and MSE. These summarize per-frame signal fidelity and enable comparison to prior work.
@@ 4 Evaluation Framework
Beyond frame-level metrics, we propose two novel perspectives for evaluation inspired by piano pedagogy: (i) an action-based evaluation (Section 4.1), in which we segment the pedal curve into press/hold/release and score temporal coverage and alignment; and (ii) a gesture-based evaluation (Section 4.2), where we consider complete pedal gestures from pedal on to off and compare characteristic contour shapes that imply pedaling in expressive piano performance.
@@ 4.1 Action-Based Evaluation
The action-based evaluation can now be framed as a three-class classification task for press, hold, release classes using action states extracted using the steps mentioned above, and the evaluation follows the standard precision, recall, and F1 framework. We report scores for each action class as well as class-balanced macro and weighted averages, all shown in Table 2.
@@ 4.2 Gesture-Based Evaluation
For all non-plain gestures in the corpus: duration (frames) has mean $147.8$, median $70.0$, std $402.9$; max-depth ratio has mean $0.645$, median $0.688$, std $0.220$. We set max-depth ratio threshold to $0.65$ and duration threshold to $100$ frames to define our quadrants.
We already see that frame-wise scores can miss contour similarity in Section 3.3. To compare gesture shapes while tolerating small local deviations, we adopt (i) Fourier-descriptor analysis: compute discrete Fourier coefficients and reconstruct with the first K=11K{=}11 coefficients to suppress high-frequency noise/model artifacts, and compare the reconstructed low-pass signals via MSE; (ii) 5-point analysis: compute duration-weighted MSE using five landmarks per gesture $x_{1:T}$: $x_{1}$, $x_{T}$, $\text{median}(x_{1:T})$, $\text{mean}(x_{1:T})$, $\text{max}(x_{1:T})$. According to these new metrics, the top example in Figure 2 yields an MSE of 0.0586 using the Fourier method, whereas the bottom example obtains an MSE of 0.0274 in the 5-point analysis, now both being less harshly graded than by raw MSE scores.
5 Experiments
5.1 Dataset and Training
[…]
[5] 0.8975 0.8971 0.8973 0.6849 0.6971 0.6863 0.0425 0.1339
Audio (Binary) 0.8946 0.8944 0.8945 0.6016 0.6627 0.6166 0.0582 0.1502
Audio 0.9043 0.9037 0.9039 0.7013 0.7153 0.7045 0.0416 0.1237
Audio+MIDI 0.9379 0.9370 0.9372 0.7529 0.7661 0.7546 0.0280 0.0986
Table 1: Frame-level results.
@@ 5.2 Evaluation Across Models
We use conventional frame-based metrics (Table 1) as well as our action- (Table 2) and gesture-based metrics (Table 3). Overall, the Audio+MIDI model performs the best, followed by Audio, with Audio (Binary) last. This pattern highlights that estimating continuous pedal values is beneficial, and adding MIDI input further enhances performance by providing helpful structural priors.
Continuous estimation matters (vs. binary classification). Although Audio (Binary) achieves a comparable frame-wise F1, it exhibits systematic bias at both action and gesture levels (Figure 3, Table 2). Despite nearly identical action distributions, Audio (Binary) achieves lower F1 scores especially for press and release actions. This gap likely arises because continuous estimation captures the smooth depth changes during transitions, whereas binary classification reduces them to abrupt on/off switches. At gesture level, Audio (Binary) largely collapses to predicting highland, and rarely recognizes more complex patterns such as mountain and pinnacle (Fig. 3). Even for highland, which Audio (Binary) predicts frequently, gesture-level MSE remains unsatisfactory (Table 3).
Models trained for continuous estimation have remarkably improved action-level recalls and modestly but consistently higher precisions than Audio (Binary) (Table 2). In addition, at the gesture level, Audio and Audio+MIDI can identify gesture categories other than highland and pinnacle, yielding distributions closer to ground truth (Figure 3), and they also result in substantially lower per-gesture MSE, a positive sign for more accurate shape modeling. These findings agree with [5] and reinforce the importance of continuous-valued prediction for piano pedal estimation.
@@ 6 Conclusion
We studied sustain pedal depth estimation with controlled experiments and evaluated using our proposed new framework with three levels at frame, action, and gesture. Across three model variants, we demonstrated that continuous-valued estimation is crucial for accurately capturing actions and preserving gesture characteristics. Pre-extracted MIDI pitch and velocity information provides useful structural priors for better contour shaping in gestures. While conventional frame-wise metrics can miss musically acceptable behavior, our proposed three-level approach provides a more comprehensive view of strengths and limitations of our models. These results show the value of our musically informed evaluation and point to future work on boundary-sensitive objectives, contour-aware losses, and perceptual validation for piano sustain pedal depth estimation tasks.